OpenStudy (anonymous):

Find the perimeter of the rectangle with length 97 inches and width 17 inches. (1 point)228 in. 211 in. 1,649 in. 114 in.

OpenStudy (anonymous):

perimeter = 2(l+W) =2(97+17) =2(114)=228

OpenStudy (anonymous):

how did you get it?

OpenStudy (anonymous):

okay thanks huh

OpenStudy (anonymous):

can you help on this one please Each pail of plaster covers 90 square feet of ceiling. What is the least number of pails of plaster you would need to buy to cover the ceiling of a room with walls 14 feet long? (1 point)5 pails 6 pails 4 pails 3 pails

OpenStudy (anonymous):

PLEASE HELP ASAP